OpenStudy (anonymous):

what is the point-slope form of a equation with the slope that passes through the point (10,-2)?

OpenStudy (gorv):

u need one more point to determine the slope

OpenStudy (anonymous):

What is the point-slope form of a equation with the slope that passes through the point (10,-2) A. y+2=3/5(x-10) B. y-2=3/5(x+10) C. y-10=3/5(x+2) D. y+10=3/5(x-2)

OpenStudy (anonymous):

A